# Topologic Attention Networks: Attending to Direct and Indirect Neighbors through Gaussian Belief Propagation

### v1

@@ Abstract
Graph Neural Networks rely on local message passing, which limits their ability to model long-range dependencies in graphs. Existing approaches extend this range through continuous-time dynamics or dense self-attention, but both suffer from high computational cost and limited scalability. We propose Topologic Attention Networks, a new framework that applies topologic attention, a probabilistic mechanism that learns how information should flow through both direct and indirect connections in a graph. Unlike conventional attention that depends on explicit pairwise interactions, topologic attention emerges from the learned information propagation of the graph, enabling unified reasoning over local and global relationships. This method achieves provides state-of-the-art performance across all measured baseline models. Our implementation is available at https://github.com/Marshall-Rosenhoover/Topologic-Attention-Networks.
@@ 1 Introduction
An intrinsic component of Graph Neural Networks (GNNs) is neighborhood aggregation: combining information from a node’s neighbors to compute its representation. Many standard GNNs, such as graph convolutional networks [GCN] and graph attention networks [GAT], rely on aggregation functions that are only determined by immediate neighbors. This local aggregation limits expressiveness and can struggle to capture long-range dependencies [struggle_denpendencies_1, struggle_denpendencies_2]. Determined to solve this problem, two main branches have emerged to aggregate information beyond the local neighborhood: continuous dynamic GNNs and graph transformers.
Continuous dynamic GNNs, such as Graph Neural ODE [GDE] and Continuous Graph Neural Networks [CGNN], extend message passing into continuous time by parameterizing information flow as a differential equation or continuous-time dynamical system. This formulation allows, in principle, unbounded propagation depth and smooth global interactions. However, these methods are prone to vanishing or exploding gradients, and remain inherently sequential [GDE, CGNN, continuous_spiking].
Graph transformers [graph_transformer, graphormer] take the opposite approach by applying self-attention over a range of node representations to capture relationships beyond immediate neighbors. By encoding structural priors into node embeddings, they can capture complex relational patterns without explicit message passing. Yet, this expressivity comes at high compute and memory costs, and the effective incorporation of graph structure remains a nontrivial challenge [transformer_nontrivial].
Figure 1: Illustration of a single head of topologic attention via a learned Gaussian Graphical Model, where node shading encodes confidence and edge color denotes interaction polarity and strength.
In this work, we propose a third branch of dynamic neighborhood aggregation: Topologic Attention Networks (TANs). Unlike continuous dynamic GNNs, which require sequential integration, or graph transformers, which depend on dense pairwise attention, TANs perform topologic attention, learning how information should flow probabilistically through the graph’s topology to capture both local and global relationships. Figure  1 illustrates a potential learned topologic attention pattern. Laying the foundation for this new branch, we present the Topologic Attention Layer, a GNN layer that implements topologic attention. We evaluate our method across diverse graph benchmarks and find that they achieve accuracy comparable to or exceeding state-of-the-art methods, highlighting the potential of topologic attention for local and global reasoning in graphs.
@@ 2 Background
A linear transformation on a graph (graph linear transformation) can be understood as a global operator that mixes node features according to the connectivity of the graph. Achieving this would require each node to integrate information from all other nodes along every possible path that connects them. However, capturing these higher-order dependencies explicitly is intractable, as enumerating all such paths is known to be NP-hard [sum_of_walks_intractable, sum_of_walks_intractable_2]. Instead of constructing this transformation explicitly, we can reinterpret the graph as a Gaussian graphical model whose equilibrium yields the solution to the graph linear transformation. In this formulation, Gaussian dependencies implicitly accumulate the influence of all paths in the graph, eliminating the need to enumerate them.
2.1 Gaussian Graphical Models
Gaussian Graphical Models represent a set of continuous random variables $x=[x_{1},\ldots,x_{N}]^{\top}$ that follow a joint multivariate Gaussian distribution with covariance matrix $\Sigma$. Each variable is associated with a node in a graph, where edges capture conditional dependencies between variables. The precision matrix $J=\Sigma^{-1}$ encodes these conditional dependencies with nonzero entries $J_{ij}$ indicate edges between nodes [malioutov2006walk, GaBP_Thesis, GaBP_intro]. Together with this structural information, the model includes an information vector $h$ representing the nodes’ local evidence. The pair $(J,h)$ fully defines the probabilistic state of the system, with the joint distribution
@@ 2.1 Gaussian Graphical Models
where $J\in\mathbb{R}^{N\times N}$, $h\in\mathbb{R}^{N}$, and $x\in\mathbb{R}^{N}$. The equilibrium of this system
$\mu=J^{-1}h$
corresponds to the solution of the graph linear transformation introduced earlier. However, directly inverting $J$ is computationally expensive and scales poorly with graph size.
2.2 Gaussian Belief Propagation
Instead of performing the matrix inversion, Gaussian Belief Propagation (GaBP) solves the system through localized message passing. In univariate GaBP, each message represents a univariate Gaussian distribution and is parameterized by two scalars: the precision $\pi$ (inverse variance) and the information $\eta$ (precision-weighted mean) [GaBP_Thesis, GaBP_WalkSum1, malioutov2006walk]. At each iteration, node $i$ sends a message to a neighbor $j$ based on incoming messages from all other neighbors except $j$, denoted $\mathcal{N}(i)\setminus j$. The message updates are defined as
$\pi_{i\to j}=-\frac{J_{ij}^{2}}{\alpha_{i\setminus j}},\qquad\eta_{i\to j}=-\frac{J_{ij}}{\alpha_{i\setminus j}}\,\beta_{i\setminus j},$
where
@@ 2.2 Gaussian Belief Propagation
After convergence, the linear system $J\mu=h$ is solved implicitly, and the resulting solution can be recovered by $\mu_{i}=\eta_{i}/\pi_{i}$.
@@
GaBP is guaranteed to converge on cyclic graphs only when the precision matrix is walk-summable. Walk-summability requires that the normalized precision matrix $\tilde{J}=D^{-1/2}JD^{-1/2}$ satisfy the spectral condition
@@
Under this condition, information propagation corresponds to a convergent geometric series,
$\sum_{k=0}^{\infty}(I-\tilde{J})^{k},$
so that contributions from longer walks decay geometrically and the total influence of all walks remains finite.
3 Topologic Attention Networks
Figure 2: Architecture of the Topologic Attention Layer.
As discussed in the previous section, constructing an explicit graph linear transformation requires enumerating all paths between nodes, which is computationally intractable. Instead, we model information propagation as probabilistic inference over a walk-summable graph topology $J$. At equilibrium, GaBP satisfies
$J\mu=h\quad\Rightarrow\quad\mu_{i}=\sum_{j}(J^{-1})_{ij}h_{j},$
where $(J^{-1})_{ij}$ quantifies how information from node $j$ influences node $i$ through all paths in the graph. In this formulation, the effective neighborhood of a node can become dynamic in two senses. First, the precision matrix $J$ determines which nodes are able to influence one another, so a node’s neighborhood is defined by the topology encoded in $J$ rather than by the raw graph adjacency. Second, the observation vector $h$ determines how strongly the nodes within this topology contribute to the aggregation: nodes with weak evidence exert little influence, while nodes with strong evidence have proportionally larger impact. Thus, $J$ defines the dynamic neighborhood, and $h$ defines the dynamic weighting within that neighborhood.
Building on this principle, we propose making the probabilistic inference learnable by parameterizing the node evidence $h$ and, optionally, the topology $J$. This allows the model to jointly learn the topological neighborhoods and how strongly each node should participate in them. We refer to this as topologic attention, implemented through the Topologic Attention Layer.
3.1 Topologic Attention Layer
@@
The overall structure of the Topologic Attention Layer, illustrated in Figure 2, consists of two subblocks: a multi-head GaBP block followed by a node-wise feed-forward network. Each subblock applies layer normalization to the node features before processing, following the pre-normalization design shown to improve training stability in Transformer architectures [prenorm]. In addition, the GaBP block applies a second layer normalization after the graph linear transformation to mitigate large, graph-dependent scale variations arising from the learned node degrees and edge weights. Since the GaBP operation itself is linear, a nonlinear activation function is applied to the updated node embeddings, which are then projected back to the original embedding dimension through a linear layer. Residual connections are incorporated around each subblock, prior to the normalization step.
3.2 Modeling Assumptions
@@
In GNNs, nodes are typically represented by multidimensional embeddings. Extending GaBP to be multivariate, where each node is multidimensional, causes the messages to be matrices that couple all embedding dimensions. To maintain scalability, we therefore make two simplifying assumptions. First, the neural network maps input features into an embedding space where dimensions can be treated as independent. Second, all dimensions share a single global precision matrix $J$. Intuitively, this means that each embedding dimension in $h$ can be treated as its own univariate GaBP and is propagated through the same topology $J$.
3.3 Precision Matrix Designs
@@
Different constructions of $J$ induce distinct inductive biases and modes of topologic attention, helping to shape whether propagation behaves as local, global, smooth, or selective. We consider three walk-summable designs of $J$: Pairwise Normal, Diagonally Dominant, and Laplacian.
@@
Pairwise-normalizable Gaussian graphical models span the same class of precision matrices as walk-summable models [malioutov2006walk]. Each edge $e=(i,j)$ contributes a local precision block
@@
This construction induces a mutual-consistency bias, since the effective strength of an edge depends jointly on both endpoints: edges where either node has low self-precision contribute weak couplings, while edges supported by two confident nodes contribute strong ones. As a result, information propagation is naturally selective, emphasizing well-supported connections and attenuating uncertain or inconsistent ones.
@@
Diagonally dominant matrices are a subset of walk-summable matrices [GaBP_Thesis]. Such matrices are characterized by having each diagonal entry larger in magnitude than the sum of the magnitudes of the off-diagonal entries in the same row, that is,
@@
In this design, each node is more confident in itself than in its connections to other nodes. This induces a confidence-centered bias, where information tends to flow outward from nodes with large diagonal precision and is absorbed more weakly by nodes with low self-precision. As a result, propagation becomes locally anchored, with high-confidence nodes acting as influential hubs and low-confidence nodes behaving primarily as receivers rather than broadcasters.
@@
Laplacian-based matrices are derived from the normalized graph Laplacian, defined as
@@
where $A$ is the weighted adjacency matrix and $D$ is the corresponding degree matrix. The eigenvalues of $L$ lie in the range $(0,2)$; by rescaling them to $(0,1)$, the resulting precision matrix is positive definite and walk-summable. In this design, information is smoothed according to the connectivity of the graph structure, inducing a manifold-smoothing bias where nodes embedded in similar structural environments co-vary strongly. As a result, propagation tends to operate at a community or cluster level, promoting broad diffusion over the graph while preserving contrasts between different structural regions.
3.4 Learnable Components of Topologic Attention
Figure 3: Depiction of how to construct a learned precision matrix $J$.
@@ 3.4 Learnable Components of Topologic Attention
The observation vector $h$ is obtained by projecting the node embeddings $X\in\mathbb{R}^{N\times d_{\text{model}}}$ into a latent observation space via a trainable transformation:
$h=\mathrm{LeakyReLU}(XW_{\text{obs}}),\qquad W_{\text{obs}}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{latent}}}.$
When the precision matrix $J$ is learned, its off-diagonal entries are derived from pairwise node similarities or reusing the observation vector. If new pairwise node similarities are used, they are computed from a learnable embedding
$s=\mathrm{LeakyReLU}(XW_{\text{sim}}),\qquad W_{\text{sim}}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{sim}}}.$
For every edge $(i,j)$, a scalar compatibility score is produced by a similarity function,
$J_{ij}=\mathrm{Similarity}(s_{i},s_{j}),$
implemented as an MLP, cosine similarity, or a Gaussian kernel. [^1] If the similarity function is not symmetric, we symmetrize the matrix as $J=(J+J^{\top})/2$. These raw scores serve as the initial edge strengths before being mapped into one of the walk-summable precision matrix constructions.
The choice of construction determines how the similarity-based edge strengths are converted into a valid precision matrix. For the learned Pairwise Normal model, cosine similarity provides the edge couplings $b_{ij}$, and nodewise self-precisions $a_{ij}$ and $c_{ij}$ are generated by neural mappings of the node embeddings and scaled to satisfy the constraint $a_{ij}c_{ij}>b_{ij}^{2}$. For the learned Diagonally Dominant model, cosine similarity parameterizes the off-diagonal terms $J_{ij}$, and the diagonal entries are formed by summing the magnitudes of the incident couplings. A learned nonnegative self-confidence term, obtained through a softplus transformation of the node embedding, allows each node to increase its own precision without strengthening its edges. For the learned Laplacian model, a Gaussian kernel on the similarity embeddings defines a weighted adjacency, which is degree-normalized and rescaled to ensure all eigenvalues lie strictly within $(0,1)$. A small learnable diagonal contribution introduces slight node-specific variation while preserving the Laplacian structure. These mappings maintain the characteristic inductive biases of each construction while allowing the topology to adapt smoothly during training.
3.5 Multi-Head GaBP
@@
The independence and parameter sharing assumptions, introduced for scalability, may be an oversimplification of the system dynamics that we want to model. To add expressiveness back into the model, we propose multi-headed GaBP, akin to multi-head attention in Transformers  [attn_all_u_need]. Through multi-headed GaBP, each head learns its own walk-summable topology $J^{(k)}$ and observation vector $h^{(k)}$, and performs GaBP independently to obtain head-specific graph linear transformations. This formulation allows different heads to focus on distinct structural, semantic, or confidence contexts within the graph. The resulting node embeddings are layer-normalized, activated, concatenated, and projected back to the feature dimension.

### v2

# Aggregating Direct and Indirect Neighbors through Graph Linear Transformations

@@ Abstract
Graph neural networks (GNN) typically rely on localized message passing, requiring increasing depth to capture long range dependencies. In this work, we introduce Graph Linear Transformations, a linear transformation that realizes direct and indirect feature mixing on graphs through a single, well-defined linear operator derived from the graph structure. By interpreting graphs as walk-summable Gaussian graphical models, we compute these transformations via Gaussian Belief Propagation, enabling each node to aggregate information from both direct and indirect neighbors without explicit enumeration of multi-hop paths. We show that different constructions of the underlying precision matrix induce distinct and interpretable propagation biases, ranging from selective edge-level interactions to uniform structural smoothing, and that Graph Linear Transformations can achieve competitive or superior performance compared to both local message-passing GNNs and dynamic neighborhood aggregation models across homophilic and heterophilic benchmark datasets.
Graph Linear Transformation, Gaussian Belief Propagation, Gaussian Graphical Models
@@ 1 Introduction
The standard idea behind feed-forward neural networks is to apply a linear transformation to the input features followed by a nonlinear activation function. Applying this concept to graphs, it would be natural to define a graph linear transformation followed by a nonlinear activation function, where the graph linear transformation corresponds to a matrix operator derived from the graph structure acting on node features. Such a transformation would aggregate information from all nodes, weighted by the structure of the graph. However, explicitly constructing such a transformation — by enumerating the contributions along all multi-hop paths in the graph — is not feasible, as this requires reasoning over an exponential number of path-dependent interactions, a problem known to be NP-hard  (Borgwardt and Kriegel, 2005; Gärtner et al., 2003).
This intractability is consistent with the practical use of localized, layer-wise approximations that propagate information through the graph in a compositional manner. In particular, Graph Convolutional Networks (Kipf and Welling, 2017) and related message-passing architectures can be viewed as approximations to graph linear transformations by repeatedly aggregating information from local neighborhoods. While effective in many settings, these methods rely on increasing network depth to approximate first order and higher order feature mixing and therefore exhibit an inherent inductive bias toward locality.
More recent research has gathered toward dynamic neighborhood aggregation in the form of attention-based models and continuous dynamic models to relax the fixed-depth limitations of local layer-wise approximations. However, these methods realize wider context through dynamic composition — either by stacking local interactions over time or by relaxing neighborhood structure to enable dense interactions — rather than through a single graph-induced linear transformation that jointly captures direct and indirect dependencies.
In this work, we ask whether graph linear transformations can be realized directly, without relying on deep stacks of local operators. We show that such transformations can be computed by imposing simple structural constraints, and that they can be either fixed or learned from data. Our results suggest that local and nonlocal graph mixing can be achieved directly, rather than through depth-based approximations.
Figure 1: Illustration of a graph being interpreted as a walk summable Gaussian graphical model, where node shading encodes confidence and edge color denotes interaction polarity and strength.
@@ 2 Background
Gaussian graphical models represent a set of continuous random variables $x=[x_{1},\ldots,x_{N}]^{\top}$ that follow a joint multivariate Gaussian distribution with covariance matrix $\Sigma$. Each variable is associated with a node in a graph, where edges capture conditional dependencies between variables. The precision matrix $J=\Sigma^{-1}$ encodes these conditional dependencies with nonzero entries $J_{ij}$ indicating edges and edge strength between nodes with $J_{ii}$ encoding the strength of a node’s self confidence  (Malioutov et al., 2006; Bickson, 2008; Ortiz et al., 2021). Together with this structural information, the model includes an information vector $h$ representing the nodes’ local evidence. The pair $(J,h)$ fully defines the probabilistic state of the system, with the joint distribution
@@
where $J\in\mathbb{R}^{N\times N}$, $h\in\mathbb{R}^{N}$, and $x\in\mathbb{R}^{N}$. The equilibrium of this system is thus given by
$\mu=J^{-1}h.$
However, directly inverting $J$ becomes computationally expensive as graph size scales. Gaussian Belief Propagation (GaBP) addresses this by solving the linear system through local message passing on the graph induced by the sparsity pattern of $J$. In the univariate setting, GaBP associates a scalar message with each directed edge $i\to j$. Each message is parameterized by a precision $\pi_{i\to j}$, current confidence in its prediction, and an information term $\eta_{i\to j}$, the current belief of the node weighted by its belief. At each iteration, node $i$ computes a message to neighbor $j$ using all incoming messages from neighbors other than $j$. Let $\mathcal{N}(i)\setminus j$ denote this set, and define
@@
The message updates are then given by
$\pi_{i\to j}=-\frac{J_{ij}^{2}}{\alpha_{i\setminus j}},\qquad\eta_{i\to j}=-\frac{J_{ij}}{\alpha_{i\setminus j}}\,\beta_{i\setminus j}.$
After convergence, the marginal parameters at each node are obtained by aggregating all incoming messages,
$\pi_{i}=J_{ii}+\sum_{k\in\mathcal{N}(i)}\pi_{k\to i},\qquad\eta_{i}=h_{i}+\sum_{k\in\mathcal{N}(i)}\eta_{k\to i},$
yielding the solution $\mu_{i}=\eta_{i}/\pi_{i}$.
@@
GaBP is guaranteed to converge on cyclic graphs only when the precision matrix is walk-summable. Under this condition, the total influence of all walks remains finite. Walk-summability, thus, requires that the normalized precision matrix $\tilde{J}=D^{-1/2}JD^{-1/2}$, where $D$ is the diagonal of the precision matrix, satisfies the spectral condition
@@
3 Graph Linear Transformations
Figure 2: Architecture of the Graph Linear Transformation Layer.
To realize graph linear transformations directly, we reinterpret graphs as walk-summable Gaussian graphical models, as illustrated in Figure 1. In this formulation, Gaussian dependencies implicitly accumulate the influence of all paths in the graph, eliminating the need to enumerate them. By parameterizing the node evidence $h$ and, optionally, the topology $J$, we introduce the graph linear transformation layer. In this, the model is able to learn how strongly each node should participate in their neighborhood and structure of the neighborhoods.
3.1 Graph Linear Transformation Layer
@@
graphical models The overall structure of the Graph Linear Transformation Layer, illustrated in Figure 2, consists of two subblocks: a Graph Linear Transformation block followed by a node-wise feed-forward network. Each subblock applies layer normalization to the node features before processing, following the pre-normalization design shown to improve training stability in Transformer architectures (Xiong et al., 2020). In addition, the Graph Linear Transformation block applies a second layer normalization after the graph linear transformation to mitigate large, graph-dependent scale variations arising from differences in node degrees and edge weights. Since the operation itself is linear, a nonlinear activation function is applied to the updated node embeddings, which are then projected back to the original embedding dimension through a linear layer. Residual connections are incorporated around each subblock, prior to the normalization step.
3.2 Creation of Graph Linear Transformations
@@
In GNNs, nodes are typically represented by multidimensional embeddings. Extending GaBP to be multivariate, where each node is multidimensional, causes the messages to be matrices that couple all embedding dimensions. To maintain scalability, we therefore make two simplifying assumptions. First, the neural network maps input features into an embedding space where dimensions can be treated as independent. Second, dimensions share a global precision matrix $J$. Intuitively, this means that each embedding dimension in $h$ can be treated as a univariate case of GaBP and is propagated through the same topology $J$.
3.2.1 Precision Matrix Designs
@@
Different constructions of $J$ induce different inductive biases of information propagation, helping to shape whether propagation behaves as local, global, smooth, or selective. We consider three walk-summable designs of $J$: Pairwise Normal, Diagonally Dominant, and Laplacian.
@@
Pairwise-normalizable Gaussian graphical models span the same class of precision matrices as walk-summable models (Malioutov et al., 2006). Each edge $e=(i,j)$ contributes a local precision block
@@
This construction induces a mutual-consistency bias, since the effective strength of an edge depends jointly on both endpoints: edges where either node has low self-confidence in the edge contribute weak couplings, while edges supported by two nodes with a high self-confidence in the edge contribute strong ones. As a result, information propagation is naturally selective, emphasizing well-supported connections and attenuating uncertain or inconsistent ones.
@@
Diagonally dominant matrices are a subset of walk-summable matrices (Bickson, 2008). Such matrices are characterized by having each diagonal entry larger in magnitude than the sum of the magnitudes of the off-diagonal entries in the same row, that is,
@@
In this design, each node is more confident in itself than in its connections to other nodes. This induces a confidence-centered bias, where information tends to flow outward from nodes with large diagonal precision and is absorbed by nodes with low self-precision. As a result, propagation becomes locally anchored, with high-confidence nodes acting as influential hubs and low-confidence nodes behaving primarily as receivers rather than broadcasters.
Figure 3: Depiction of how to construct a precision matrix $J$.
@@
Laplacian-based matrices are not inherently walk-summable matrices as their eigenvalues inherently lie in the range [1,1][-1,1]. They are derived from the normalized graph Laplacian as
@@
where $A$ is the weighted adjacency matrix and $D$ is the corresponding degree matrix. By rescaling $D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$ to lie in the range (1,1)(-1,1), the resulting matrix becomes walk-summable. In this, information is smoothed according to the connectivity of the graph structure, inducing a manifold-smoothing bias where nodes embedded in similar structural environments co-vary strongly. As a result, propagation tends to operate at a community or cluster level, promoting broad diffusion over the graph while preserving contrasts between different structural regions.
3.2.2 Parameterization of Graph Linear Transformations
@@
For the learnable components of the graph linear transformation, the observation vector $h$ and precision matrix $J$ are parameterized from base node embeddings. The observation vector $h$ is obtained by projecting the node embeddings $X\in\mathbb{R}^{N\times d_{\text{model}}}$ into a latent observation space,
$h=\text{LeakyReLU}(XW_{\text{obs}}),\qquad W_{\text{obs}}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{latent}}}.$
To construct the precision matrix $J$, edge-level interactions are derived either from a fixed graph structure or from learned pairwise node similarities. The overall construction process is summarized in Figure 3. In the learned setting, node embeddings are first mapped into a similarity space, either by directly using the observation vector $h$ to compute similarity-based edge strengths or by learning a separate similarity projection,
$s=\text{LeakyReLU}(XW_{\text{sim}}),\qquad W_{\text{sim}}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{sim}}},$
after which a scalar compatibility score is computed with a similarity function  [^1] Since the precision matrix must be symmetric, non-symmetric similarity functions require an explicit symmetrization step, such as $(J+J^{T})/2$. for each edge $(i,j)$,
$J_{ij}=\text{Similarity}(s_{i},s_{j}).$
The resulting interaction strengths are subsequently mapped into one of the walk-summable precision matrix constructions introduced in Section 3.2.1. The exact implementation of the learned and fixed precision matrices used in this paper in detail in Appendix section A.1.
3.2.3 Multi-Head Graph Linear Transformations
@@
The independence and parameter sharing assumptions, introduced for scalability, may be an oversimplification of the system dynamics that we want to model. To add expressiveness back into the model, we propose multi-headed graph linear transformations for learned precision matrices, akin to multi-head attention in Transformers  (Vaswani et al., 2017). Through multi-headed Graph Linear Transformations, each head learns its own walk-summable topology $J^{(k)}$ and observation vector $h^{(k)}$, and performs a graph linear transformation independently to obtain head-specific graph linear transformation results. This formulation allows different heads to focus on distinct structural, semantic, or confidence contexts within the graph. The resulting node embeddings are layer-normalized, activated, concatenated, and projected back to the feature dimension.